An automobile company has extra capacity that can be used to produce gears that the company has been buying for Rs. 300 each. If the company makes gears, it will incur materials cost of Rs. 90 per unit, labour costs of Rs. 120 per unit, and variable overhead costs of Rs. 30 per unit. The annual fixed cost associated with the unused capacity is Rs. 2,40,000. Demand for next year is estimated at4,000 units

.(a)Would it be profitable for the company to make the gears?

(b)Suppose the capacity could be used by another department for the production of some agricultural equipment that would cover its fixed and variable cost and contribute Rs. 90,000 to profit which would be more advantageous, gear production or agricultural equipment production?
Answer #1

a) Current cost of buying the gears from outside = Rs. 300 * 4,000 = Rs. 1,200,000

Cost of producing gears using excess capacity for each unit = Material + Labour + Variable overhead = Rs. 90 + 120 + 30 = Rs. 240

(fixed cost is not taken above since it is being incurred anyway)

Total cost of inhouse production of gears = Rs. 240 * 400,000 = Rs. 960,000

Net benefit to the profit = Rs. 1,200,000 - 960,000 = Rs. 240,000

b) Assuming that "... would cover its fixed and variable cost" means that the Rs. 240,000 which is the current fixed cost of this unused capacity will be recovered. So the net benefit woudl be Rs. 240,000 + 90,000 = Rs. 330,000

It would be more beneficial to produce agricultural equipment.